OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 28 feet, what is its lateral area? use pi= 22/7

OpenStudy (anonymous):

if area = 154, then radius = \[\sqrt{154/\pi}\] = 7 feet. then circumference = 2 pi r = 44 feet then lateral area = 44 * 28 = 1232 feet.

OpenStudy (anonymous):

well, square feet :P

OpenStudy (anonymous):

ok, so I had to find out the area formula of the cylinder first. Then reverse it.

OpenStudy (anonymous):

nononononon, just find the area formula of the base, then reverse it to find the radius, and then find the circumference. the circumference * lenght of the cylinder will give you the curved, or lateral area.

OpenStudy (anonymous):

right, I just didn't know how to do the first part. Thank you :)

OpenStudy (anonymous):

no problem :D